# Energy-Enhanced Expansion of the Standard Model Effective Field Theory

### v1

@@ Abstract
We formalize energy-scaling arguments in the Standard Model Effective Field Theory (SMEFT) to estimate effects of operators up to dimension ten. Introducing a classification based on the number of external legs and an energy-counting parameter, we establish a dual expansion in $v/\Lambda$ and $E/\Lambda$. Extending to four-, five-, and six-particle vertices, our framework highlights energy-enhanced operators that dominate high-energy processes at the HL-LHC. This organization streamlines experimental analyses to only include operators with energetic impact in their analyses and enhances the discoverability of new physics within the SMEFT framework.
@@ I Introduction
From the perspective of model building and theoretical investigations, our classification highlights the characteristics of new physics models that render them more accessible to discovery at the HL-LHC. Specifically, models that generate energy-enhanced operators, particularly those arising from tree-level matching conditions, tend to be more discoverable within this framework. Importantly, these criteria do not alter the SMEFT in any way but rather provide a systematic organizational framework. Moreover, we offer a systematic way to manage the rapid proliferation of operators that arise at higher mass dimensions Henning:2015alf . This focused approach is especially beneficial in precision phenomenology, where combining data from different energy scales necessitates careful renormalization group (RG) evolution Assi:2023zid ; Chala:2021pll ; DasBakshi:2022mwk ; Bakshi:2024wzz ; Boughezal:2024zqa ; Helset:2022pde . In effect, by concentrating on the most promising operators, the overall theoretical framework becomes more tractable, allowing for a clearer connection between experimental observations and the underlying EFT. Indeed, our vector‑boson fusion (VBF) analysis at the HL‑LHC Assi:2024zap is a prime example of how this discoverability‑focused strategy can enhance sensitivity and provide a clear link between experimental observables and the underlying EFT contributions. While Ref. Assi:2024zap focused on VBF Higgs production, other studies have successfully applied similar ideas to different processes Kim:2022amu ; Martin:2023tvi ; Corbett:2023yhk ; Degrande:2023iob .
@@ II Energy Expansion
where $i$ indexes the mass dimension above four, and $c_{j}^{(i)}$ and $\mathcal{O}_{j}^{(4+i)}$ represent the Wilson coefficients and corresponding operators of dimension $4+i$, respectively. In the following we assume that $c_{j}^{(i)}/\Lambda^{i}$ are all of comparable magnitude—an assumption that reflects a weakly coupled new physics scenario. In the absence of a specific UV model that enforces a strong hierarchy (for example, via a symmetry or dynamical mechanism), there is no a priori reason to expect that these coefficients should differ substantially. This is reasonable to assume since generically, matching calculations in perturbative UV completions tend to produce contributions of similar size across different operators deBlas:2014mba ; deBlas:2017xtg ; Carmona:2021xtq ; Fuentes-Martin:2022jrf ; Li:2023cwy . We will discuss later how deviations from this assumption affect the power counting.
The Lagrangian in Eq. (2) post Electroweak Symmetry Breaking (EWSB) can be ordered by the number of particles in a tree-level vertex. This is valid since not all operators will contribute to a vertex with a given number of legs; operators containing many fields won’t contribute at tree level to vertices with few legs, while operators with only a few fields can’t contribute to vertices with many legs. Thus, for operators that can contribute to a given vertex leg multiplicity, we can derive how each scales with $v$ and $E$, the two scales in the problem.
@@
The organization in Eq. (3) extends this energy classification concept to higher multiplicity ($\geq 4$ particle) vertices. As the kinematics is no longer trivial (momentum products do not just reduce to combinations of masses), the $E$ scaling is not set by the lowest dimension operator contributing to a given vertex. While this means more complicated scalings are possible, and the scaling can change as we increase the mass dimension, it is straightforward to determine the scaling, at as we will show for dimension six through to dimension ten operators. This suffices for the purposes of identifying which effects through $\mathcal{O}(1/\Lambda^{6})$ are the most likely to impact tails of kinematic distributions – assuming that all Wilson coefficients are the same order of magnitude.
@@
In contrast, for vertices with $n<4$ the situation is fundamentally different. With only two or three external legs, momentum conservation and on‐shell conditions fully constrain the kinematics, leaving no free invariant to carry an independent factor of $E$. In these cases, derivative factors are either fully fixed by the particle masses or can be reshuffled using equations of motion and integration‐by‐parts, so the naive $E/\Lambda$ counting does not necessarily reflect the actual energy scaling. Thus, our counting method is only robust for $n\geq 4$ because the extra independent kinematic degrees of freedom guarantee that energy enhancements appear as expected, whereas for $n<4$ the tight kinematic constraints limit the validity of an $E/\Lambda$ expansion. For $2$‑ and $3$‑point vertices, we explicitly correct for these limitations, ensuring that the energy scaling tables (see Tables 2 and 3) can be used without concern.
@@ III Further Enhancements
Second, the chirality structure of an operator is important; operators with mixed chirality (e.g. $e_{c}L,\,u_{c}Q$) tend to be suppressed if flavor universality or minimal flavor violation is assumed DAmbrosio:2002vsn ; Brivio:2017vri . To clearly convey these aspects in our summary tables, we have adopted a color–coding scheme:
  • Operators with a tree–level origin are marked in black.
  • Operators generated only at loop level are marked in orange.
  • Separately, if an operator has mixed chirality, it is flagged in blue.
  • When both features apply—for example, a loop–induced operator with mixed chirality structure—the labels are combined and such operators are shown in purple.
We employ this coding throughout the remaining sections.
@@ IV Energy Scaling Tables
Below we summarize the energy–scaling behavior for SMEFT operators contributing to vertices with varying numbers of external legs. We also incorporate the color–coding scheme introduced in Section III.
@@
${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}H^{2}X^{2}}$ $\lambda^{4}$ $(\partial V)^{2}:{v^{2}}$
${\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}\psi^{2}H^{3}}$ $\lambda^{4}$ $\psi^{2}:{v^{2}}$
@@
${\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}\psi^{2}H^{5}}$ $\lambda^{8}$ $\psi^{2}:{v^{4}}$
@@
${\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}\psi^{2}H^{7}}$ $\lambda^{12}$ $\psi^{2}:{v^{6}}$
@@
${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}H^{2}X^{2}}$ $\lambda$ $hVV:\displaystyle{v\,E^{2}}$
${\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}\psi^{2}H^{3}}$ $\lambda^{2}$ $h\,\psi^{2}:\displaystyle{v^{2}E}$
${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}X^{3}}$ $1$ $VVV:\displaystyle{E^{3}}$
${\color[rgb]{0.55859375,0.140625,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.55859375,0.140625,0.66796875}\psi^{2}HX}$ $\lambda$ $\psi^{2}\,V:\displaystyle{v\,E^{2}}$
@@
${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}H^{2}X^{3}}$ $\lambda^{4}$ $VVV:\displaystyle{v^{2}E^{3}}$
@@
${\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}\psi^{2}H^{3}X}$ $\lambda^{5}$ $\psi^{2}\,V:\displaystyle{v^{3}E^{2}}$
${\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}\psi^{2}H^{5}}$ $\lambda^{6}$ $h\,\psi^{2}:\displaystyle{v^{4}E}$
@@
${\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}\psi^{2}H^{5}X}$ $\lambda^{9}$ $\psi^{2}\,V:\displaystyle{v^{5}E^{2}}$
${\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}\psi^{2}H^{7}}$ $\lambda^{10}$ $h\,\psi^{2}:\displaystyle{v^{6}E}$
@@
${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}H^{2}X^{2}}$ $\lambda^{2}$ $h^{2}V^{2}:{E^{2}},\;\;hV^{3}:{vE},\;\;V^{4}:{v^{2}}$
${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}X^{3}}$ $\lambda^{2}$ $V^{4}:{E^{2}}$
${\color[rgb]{0.55859375,0.140625,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.55859375,0.140625,0.66796875}\psi^{2}HX}$ $\lambda^{2}$ $\psi^{2}V^{2}:{vE},\;\;\psi^{2}\,h\,\partial V:{E^{2}}$
@@
${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}X^{4}}$ $\lambda^{4}$ $V^{4}:{E^{4}}$
@@
${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}H^{2}X^{2}D^{2}}$ $\lambda^{4}$ $(\partial h)^{2}V^{2}:{E^{4}},\;\;(\partial h)V^{3}:{E^{3}v},\;\;V^{4}:{E^{2}v^{2}}\;\;$
@@
${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}\psi^{2}X^{2}D}$ $\lambda^{4}$ $\psi^{2}(\partial^{2}V)(\partial V):{E^{4}}$
${\color[rgb]{0.55859375,0.140625,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.55859375,0.140625,0.66796875}\psi^{2}HXD^{2}}$ $\lambda^{4}$ $\psi^{2}(\partial h)(\partial V):{E^{4}},\;\;\psi^{2}(\partial V)^{2}:{E^{3}v}\;\;$
@@
$H^{4}D^{6}$ $\lambda^{6}$ $(\partial h)^{4}:{E^{6}},\;\;h^{3}V:{E^{5}v},\;\;h^{2}V^{2}:{E^{4}v^{2}},\;\;hV^{3}:{E^{3}\,v^{3}},\;\;V^{4}:{E^{4}\,v^{2}}\;\;$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X^{4}D^{2}}$ $\lambda^{6}$ $(\partial V)^{2}V^{2}:{E^{6}}\;\;$
${\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}\psi^{2}HXD^{4}}$ $\lambda^{6}$ $\psi^{2}(\partial h)(\partial V):{E^{6}},\;\;\psi^{2}(\partial V)^{2}:{E^{5}v}\;\;$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\psi^{2}H^{2}D^{5}}$ $\lambda^{6}$ $(\partial h)^{2}(\partial\psi)^{2}:{E^{6}},\;\;(\partial h)(\partial\psi)^{2}V:{E^{5}v},\;\;(\partial\psi)^{2}V^{2}:{E^{4}v^{2}}\;\;$
@@
${\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}\psi^{2}H^{3}}$ $\lambda^{4}$ $\psi^{2}h^{3}:{E}\;\;$
${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}H^{2}X^{2}}$ $\lambda^{4}$ $h^{2}(\partial V)V^{2}:{E},\;\;hV^{4}:{v}\;\;$
${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}X^{3}}$ $\lambda^{4}$ $(\partial V)V^{4}:{E}$
${\color[rgb]{0.55859375,0.140625,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.55859375,0.140625,0.66796875}\psi^{2}HX}$ $\lambda^{4}$ $\psi^{2}hV^{2}:{E}\;\;$
@@
${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}X^{4}}$ $\lambda^{6}$ $(\partial V)^{3}V^{2}:{E^{3}}$
@@
${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}H^{2}X^{3}}$ $\lambda^{6}$ $h^{2}(\partial V)^{3}:{E^{3}},\;\;h(\partial V)^{2}V^{2}:{E^{2}v},\;\;(\partial V)V^{4}:{Ev^{2}}\;\;$
${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}H^{2}X^{2}D^{2}}$ $\lambda^{6}$ $(\partial h)^{2}(\partial V)V^{2}:{E^{3}},\;\;(\partial h)(\partial V)^{2}V^{2}:{E^{2}v},\;\;(\partial V)^{3}V^{2}:{Ev^{2}}\;\;$
@@
${\color[rgb]{0.55859375,0.140625,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.55859375,0.140625,0.66796875}\psi^{2}HX^{2}}$ $\lambda^{6}$ $\psi^{2}h(\partial V)^{2}:{E^{3}},\ ;\;\psi^{2}(\partial V)V^{2}:{E^{2}v}\;\;$
@@
${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}\psi^{2}X^{2}D}$ $\lambda^{6}$ $\psi^{2}(\partial V)^{2}V:{E^{3}}$
${\color[rgb]{0.55859375,0.140625,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.55859375,0.140625,0.66796875}\psi^{2}HXD^{2}}$ $\lambda^{6}$ $\psi^{2}(\partial h)(\partial V):{E^{3}},\;\;\psi^{2}(\partial V)^{2}:{E^{2}v}\;\;$
${\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}\psi^{2}H^{3}D^{2}}$ $\lambda^{6}$ $\psi^{2}(\partial h)^{3}:{E^{3}},\;\;\psi^{2}(\partial h)^{2}V:{E^{2}v},\;\;\psi^{2}(\partial h)V^{2}:{Ev^{2}}\;\;$
@@
${\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}\psi^{4}HD}$ $\lambda^{6}$ $\psi^{4}(\partial h):{E^{3}},\;\;\psi^{4}V:{E^{2}v}\;\;$
@@
$H^{4}D^{6}$ $\lambda^{8}$ $(\partial h)^{4}(\partial^{2}V):{E^{5}},\;\;(\partial h)^{3}(\partial V)^{2}:{E^{4}v},\;\;(\partial h)^{2}(\partial V)V^{2}:{E^{3}v^{2}},\;\;(\partial h)V^{4}:{E^{2}\,v^{3}},\;\;V^{5}:{E\,v^{2}}\;\;$
@@
${\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}\psi^{2}HXD^{4}}$ $\lambda^{8}$ $\psi^{2}(\partial^{2}h)(\partial V)(\partial V):{E^{5}},\;\;\psi^{2}(\partial V)^{3}:{E^{4}v}$
${\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}\psi^{2}H^{3}D^{4}}$ $\lambda^{8}$ $\psi(\partial\psi)(\partial h)^{3}:{E^{5}},\;\;\psi(\partial\psi)(\partial h)^{2}V:{E^{4}v},\;\;\psi(\partial\psi)(\partial h)V^{2}:{E^{3}v^{2}},\;\;\psi(\partial\psi)V^{3}:{E^{2}v^{3}}$
${\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}\psi^{2}HX^{2}D^{2}}$ $\lambda^{8}$ $\psi^{2}(\partial h)(\partial^{2}V)V:{E^{5}},\;\;\psi^{2}(\partial V)V^{2}:{E^{4}v}$
@@
$\psi^{4}D^{4}$ $\lambda^{8}$ $(\psi\partial\psi)(\psi\partial^{2}\psi)\partial V:{E^{5}}$
@@
${\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}\psi^{4}HD^{3}}$ $\lambda^{8}$ $(\psi\partial\psi)(\psi\partial\psi)(\partial h):{E^{5}},\;\;(\psi\partial\psi)(\psi\partial\psi)V:{E^{4}v}$

### v2

@@ Abstract
We formalize energy-scaling arguments in the Standard Model Effective Field Theory (SMEFT) to estimate effects of operators up to dimension ten. Our approach relies on weakly-coupled UV completions with no presumed large hierarchies between Wilson coefficients. We introduce a classification based on the number of external legs and an energy-counting parameter, we establish a dual expansion in $v/\Lambda$ and $E/\Lambda$. Extending to four-, five-, and six-particle vertices, our framework highlights energy-enhanced operators that dominate high-energy processes at the HL-LHC. This organization streamlines experimental analyses to only include operators with energetic impact in their analyses and enhances the discoverability of new physics within the SMEFT framework.
@@ I Introduction
From the perspective of model building and theoretical investigations, our classification highlights the characteristics of new physics models that render them more accessible to discovery at the HL-LHC. Specifically, models that generate energy-enhanced operators, particularly those arising from tree-level matching conditions, tend to be more discoverable within this framework. One can then combine our operator counting with explicit vertex-generation automated tools such as Refs. Dedes:2017zog ; Dedes:2023zws , to immediately translate into full amplitudes ready for experimental fits. Importantly, these criteria do not alter the SMEFT in any way but rather provide a systematic organizational framework.
Moreover, we offer a systematic way to manage the rapid proliferation of operators that arise at higher mass dimensions Henning:2015alf . This focused approach is especially beneficial in precision phenomenology, where combining data from different energy scales necessitates careful renormalization group (RG) evolution Assi:2023zid ; Chala:2021pll ; DasBakshi:2022mwk ; Bakshi:2024wzz ; Boughezal:2024zqa ; Helset:2022pde . In effect, by concentrating on the most promising operators, the overall theoretical framework becomes more tractable, allowing for a clearer connection between experimental observations and the underlying EFT. Indeed, our vector‑boson fusion (VBF) analysis at the HL‑LHC Assi:2024zap is a prime example of how this discoverability‑focused strategy can enhance sensitivity and provide a clear link between experimental observables and the underlying EFT contributions. While Ref. Assi:2024zap focused on VBF Higgs production, other studies have successfully applied similar ideas to different processes Kim:2022amu ; Martin:2023tvi ; Corbett:2023yhk ; Degrande:2023iob .
@@ II Energy Expansion
[…]
The Lagrangian in Eq. (2) post Electroweak Symmetry Breaking (EWSB) can be ordered by the number of particles in a tree-level vertex. This is valid since not all operators will contribute to a vertex with a given number of legs; operators containing many fields will not contribute at tree level to vertices with few legs, while operators with only a few fields cannot contribute to vertices with many legs. Thus, for operators that can contribute to a given vertex leg multiplicity, we can derive how each scales with $v$ and $E$, the two scales in the problem.
@@
The organization in Eq. (3) extends this energy classification concept to higher multiplicity (4$\geq$ 4≥ 4 particle) vertices. As the kinematics is no longer trivial (momentum products do not just reduce to combinations of masses), the $E$ scaling is not set by the lowest dimension operator contributing to a given vertex. While this means more complicated scalings are possible, and the scaling can change as we increase the mass dimension, it is straightforward to determine the scaling, at as we will show for dimension six through to dimension ten operators. This suffices for the purposes of identifying which effects through $\mathcal{O}(1/\Lambda^{6})$ are the most likely to impact tails of kinematic distributions – assuming that all Wilson coefficients are the same order of magnitude.
@@
In contrast, for vertices with $n<4$ the situation is fundamentally different. With only two or three external legs, momentum conservation and on‐shell conditions fully constrain the kinematics, leaving no free invariant to carry an independent factor of $E$. In these cases, derivative factors are either fully fixed by the particle masses or can be reshuffled using equations of motion and integration‐by‐parts, so the naive $E/\Lambda$ counting does not necessarily reflect the actual energy scaling. Thus, our counting method is only robust for $n\geq 4$ because the extra independent kinematic degrees of freedom guarantee that energy enhancements appear as expected, whereas for $n<4$ the tight kinematic constraints limit the validity of an $E/\Lambda$ expansion. For 222‑ and 333‑point vertices, we explicitly correct for these limitations, ensuring that the energy scaling tables (see Tables 2 and 3) can be used without concern.
@@ III Further Enhancements
Second, the chirality structure of an operator is important; operators with mixed chirality (e.g. $e_{c}L,\,u_{c}Q$) tend to be suppressed if flavor universality or minimal flavor violation is assumed DAmbrosio:2002vsn ; Brivio:2017vri . To clearly convey these aspects in our summary tables, we have adopted a coding scheme:
  • Operators with a tree-level origin are typeset plainly.
  • Operators generated only at loop level are typeset in bold.
  • Operators with mixed chirality are underlined.
  • Operators with both features—e.g., loop-induced and mixed chirality—are bold and underlined.
We employ this coding throughout the remaining sections. We further note that the $\lambda$ weights are helicity-blind since no assumption is made about whether the vertex leg ends up external or internal; once external $V_{L}$ or $V_{T}$ polarisations are fixed for a specific process, one should retain the operator structures that couple to those helicities. For a specific process example, see the energy-enhanced VBF analysis of Ref. Assi:2024zap .
@@ IV Energy Scaling Tables
Below we summarize the energy–scaling behavior for SMEFT operators contributing to vertices with varying numbers of external legs. We also incorporate the coding scheme introduced in Section III.
@@
$\bm{H^{2}X^{2}}$ $\lambda^{4}$ $(\partial V)^{2}:{v^{2}}$
$\underline{\vphantom{y}\psi^{2}H^{3}}$ $\lambda^{4}$ $\psi^{2}:{v^{2}}$
@@
$\underline{\vphantom{y}\psi^{2}H^{5}}$ $\lambda^{8}$ $\psi^{2}:{v^{4}}$
@@
$\underline{\vphantom{y}\psi^{2}H^{7}}$ $\lambda^{12}$ $\psi^{2}:{v^{6}}$
@@
$\bm{H^{2}X^{2}}$ $\lambda$ $hVV:\displaystyle{v\,E^{2}}$
$\underline{\vphantom{y}\psi^{2}H^{3}}$ $\lambda^{2}$ $h\,\psi^{2}:\displaystyle{v^{2}E}$
$\bm{X^{3}}$ 111 $VVV:\displaystyle{E^{3}}$
$\underline{\vphantom{y}\bm{\psi^{2}HX}}$ $\lambda$ $\psi^{2}\,V:\displaystyle{v\,E^{2}}$
@@
$\bm{H^{2}X^{3}}$ $\lambda^{4}$ $VVV:\displaystyle{v^{2}E^{3}}$
@@
$\underline{\vphantom{y}\psi^{2}H^{3}X}$ $\lambda^{5}$ $\psi^{2}\,V:\displaystyle{v^{3}E^{2}}$
$\underline{\vphantom{y}\psi^{2}H^{5}}$ $\lambda^{6}$ $h\,\psi^{2}:\displaystyle{v^{4}E}$
@@
$\underline{\vphantom{y}\psi^{2}H^{5}X}$ $\lambda^{9}$ $\psi^{2}\,V:\displaystyle{v^{5}E^{2}}$
$\underline{\vphantom{y}\psi^{2}H^{7}}$ $\lambda^{10}$ $h\,\psi^{2}:\displaystyle{v^{6}E}$
@@
$\bm{H^{2}X^{2}}$ $\lambda^{2}$ $h^{2}V^{2}:{E^{2}},\;\;hV^{3}:{vE},\;\;V^{4}:{v^{2}}$
$\bm{X^{3}}$ $\lambda^{2}$ $V^{4}:{E^{2}}$
$\underline{\vphantom{y}\bm{\psi^{2}HX}}$ $\lambda^{2}$ $\psi^{2}V^{2}:{vE},\;\;\psi^{2}\,h\,\partial V:{E^{2}}$
@@
$\bm{X^{4}}$ $\lambda^{4}$ $V^{4}:{E^{4}}$
@@
$\bm{H^{2}X^{2}D^{2}}$ $\lambda^{4}$ $(\partial h)^{2}V^{2}:{E^{4}},\;\;(\partial h)V^{3}:{E^{3}v},\;\;V^{4}:{E^{2}v^{2}}\;\;$
@@
$\bm{\psi^{2}X^{2}D}$ $\lambda^{4}$ $\psi^{2}(\partial^{2}V)(\partial V):{E^{4}}$
$\underline{\vphantom{y}\bm{\psi^{2}HXD^{2}}}$ $\lambda^{4}$ $\psi^{2}(\partial h)(\partial V):{E^{4}},\;\;\psi^{2}(\partial V)^{2}:{E^{3}v}\;\;$
@@
$H^{4}D^{6}$ $\lambda^{6}$ $\partial^{2}(\partial h)^{4}:{E^{6}},\;\;(\partial h)^{3}\partial^{2}V:{E^{5}v},\;\;(\partial h)^{2}(\partial V)V:{E^{4}v^{2}},\;\;(\partial^{3}h)V^{3}:{E^{3}\,v^{3}},\;\;\partial^{2}V^{4}:{E^{4}\,v^{2}}\;\;$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X^{4}D^{2}}$ $\lambda^{6}$ $\partial^{2}(\partial V)^{4}:{E^{6}}\;\;$
$\underline{\vphantom{y}\psi^{2}HXD^{4}}$ $\lambda^{6}$ $\partial^{2}\psi^{2}(\partial^{2}h)(\partial V):{E^{6}},\;\;\partial^{2}\psi^{2}(\partial V)^{2}:{E^{5}v}\;\;$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\psi^{2}H^{2}D^{5}}$ $\lambda^{6}$ $\partial(\partial h)^{2}(\partial\psi)^{2}:{E^{6}},\;\;(\partial h)(\partial\psi)^{2}\partial V:{E^{5}v},\;\;(\partial\psi)^{2}\partial V^{2}:{E^{4}v^{2}}\;\;$
@@
$\underline{\vphantom{y}\psi^{2}H^{3}}$ $\lambda^{4}$ $\psi^{2}h^{3}:{E}\;\;$
$\bm{H^{2}X^{2}}$ $\lambda^{4}$ $h^{2}(\partial V)V^{2}:{E},\;\;hV^{4}:{v}\;\;$
$\bm{X^{3}}$ $\lambda^{4}$ $(\partial V)V^{4}:{E}$
$\underline{\vphantom{y}\bm{\psi^{2}HX}}$ $\lambda^{4}$ $\psi^{2}hV^{2}:{E}\;\;$
@@
$\bm{X^{4}}$ $\lambda^{6}$ $(\partial V)^{3}V^{2}:{E^{3}}$
@@
$\bm{H^{2}X^{3}}$ $\lambda^{6}$ $h^{2}(\partial V)^{3}:{E^{3}},\;\;h(\partial V)^{2}V^{2}:{E^{2}v},\;\;(\partial V)V^{4}:{Ev^{2}}\;\;$
$\bm{H^{2}X^{2}D^{2}}$ $\lambda^{6}$ $(\partial h)^{2}(\partial V)V^{2}:{E^{3}},\;\;(\partial h)(\partial V)^{2}V^{2}:{E^{2}v},\;\;(\partial V)^{3}V^{2}:{Ev^{2}}\;\;$
@@
$\underline{\vphantom{y}\bm{\psi^{2}HX^{2}}}$ $\lambda^{6}$ $\psi^{2}h(\partial V)^{2}:{E^{3}},\ ;\;\psi^{2}(\partial V)V^{2}:{E^{2}v}\;\;$
@@
$\bm{\psi^{2}X^{2}D}$ $\lambda^{6}$ $\psi^{2}(\partial V)^{2}V:{E^{3}}$
$\underline{\vphantom{y}\bm{\psi^{2}HXD^{2}}}$ $\lambda^{6}$ $\psi^{2}(\partial h)(\partial V):{E^{3}},\;\;\psi^{2}(\partial V)^{2}:{E^{2}v}\;\;$
$\underline{\vphantom{y}\psi^{2}H^{3}D^{2}}$ $\lambda^{6}$ $\psi^{2}(\partial h)^{3}:{E^{3}},\;\;\psi^{2}(\partial h)^{2}V:{E^{2}v},\;\;\psi^{2}(\partial h)V^{2}:{Ev^{2}}\;\;$
@@
$\underline{\vphantom{y}\psi^{4}HD}$ $\lambda^{6}$ $\psi^{4}(\partial h):{E^{3}},\;\;\psi^{4}V:{E^{2}v}\;\;$
@@
$H^{4}D^{6}$ $\lambda^{8}$ $(\partial h)^{4}(\partial V):{E^{5}},\;\;(\partial h)^{3}(\partial V)V:{E^{4}v},\;\;(\partial h)^{2}(\partial V)V^{2}:{E^{3}v^{2}},\;\;(\partial h)V^{5}:{E^{2}\,v^{3}},\;\;(\partial V)V^{4}:{E\,v^{2}}\;\;$
@@
$\underline{\vphantom{y}\psi^{2}HXD^{4}}$ $\lambda^{8}$ $\psi^{2}(\partial^{2}h)(\partial V)(\partial V):{E^{5}},\;\;\psi^{2}(\partial V)^{3}:{E^{4}v}$
$\underline{\vphantom{y}\psi^{2}H^{3}D^{4}}$ $\lambda^{8}$ $\psi(\partial\psi)(\partial h)^{3}:{E^{5}},\;\;\psi(\partial\psi)(\partial h)^{2}V:{E^{4}v},\;\;\psi(\partial\psi)(\partial h)V^{2}:{E^{3}v^{2}},\;\;\psi(\partial\psi)V^{3}:{E^{2}v^{3}}$
$\underline{\vphantom{y}\psi^{2}HX^{2}D^{2}}$ $\lambda^{8}$ $\psi^{2}(\partial h)(\partial^{2}V)V:{E^{5}},\;\;\psi^{2}(\partial V)V^{2}:{E^{4}v}$
@@
$\psi^{4}D^{4}$ $\lambda^{8}$ $(\psi\partial\psi)(\psi\partial^{2}\psi)V:{E^{5}}$
@@
$\underline{\vphantom{y}\psi^{4}HD^{3}}$ $\lambda^{8}$ $(\psi\partial\psi)(\psi\partial\psi)(\partial h):{E^{5}},\;\;(\psi\partial\psi)(\psi\partial\psi)V:{E^{4}v}$